# A Mess of Memory System Benchmarking, Simulation and Application Profiling

### v1

@@ III Mess simulator
In this section we will present the Mess memory simulator and show how it significantly improves the memory simulation accuracy.
III-A Design
@@
The CPU and memory simulators are typically connected in the following way: the CPU simulator issues the memory operations and the memory simulator determines their latencies. The Mess simulator does this based on the application’s position in the memory bandwidth–latency curves. This process is complex due to the inherent dependency between the memory system latency, timings of the memory operations and all dependent instructions, and the generated memory bandwidth. We simplify the problem by designing the Mess simulator not to compute the exact memory latency for a given memory traffic, but to detect and correct discrepancies between the memory access latency and the simulated bandwidth. This approach, together with the fundamental principle of application’s position in the memory bandwidth–latency curves, enables the Mess simulator to surpass the accuracy of all other memory simulators, while remaining simple and fast.
@@
The Mess simulator acts as a feedback controller [77] from the classical control theory [78], illustrated in Figure 7. The CPU simulation can start from any memory access latency, e.g. the unloaded one. The feedback control loop monitors the simulated bandwidth and controls whether it correspond to the memory latency used in the CPU simulation. If this is not the case, the memory latency is being adjusted with an iterative process we describe later. The control process is performed at the end of each simulation window, which, in our experiments, comprises 1000 memory operations. This is much smaller than the length of the application phases [64], so the error due to the transition between different application phases has a negligible impact on the simulator’s accuracy.
Figure 8 describes one iteration of the Mess simulator control loop. We start with the Mess estimate of the application’s bandwidth–latency position in the $i^{th}$ simulation window, $(messBW_{i},~{}Latency_{i})$1. From that point on, all the issued memory requests are simulated with $Latency_{i}$. At the end of the simulation window, based on the outcomes of the CPU simulation, the Mess simulator monitors the simulated memory bandwidth, $cpuBW_{i}$2, and compares it with the $messBW_{i}$ estimated at the beginning of the window  3.
If the simulation is in a steady-state and the application did not change its behavior, there will be no major difference between $cpuBW_{i}$ and $messBW_{i}$. This confirms the consistency in the simulated memory access latency, the CPU timings and the achieved memory bandwidth. Therefore, the CPU simulation in the next window will continue with the same memory latency: $Latency_{i+1}=Latency_{i}$.
Otherwise, a difference between $cpuBW_{i}$ and $messBW_{i}$ suggests inconsistent simulated memory latency and bandwidth. This can happen, for example, if the application changes its behavior. Figure 8 illustrates the case in which the application increases the frequency of memory request leading to the higher bandwidth: $cpuBW_{i}>messBW_{i}$3. In this case, the simulated memory bandwidth $cpuBW_{i}$ does not correspond to the memory $Latency_{i}$ used in the CPU simulation. To address this inconsistency, the Mess simulator adjusts the predicted application position in the bandwidth–latency curves. The objective of this adjustment is not to reach the correct $(BW,Latency)$ position in a single iteration. The next Mess estimate, $(messBW_{i+1},~{}Latency_{i+1})$, will be positioned in-between $messBW_{i}$ and $cpuBW_{i}$4. The exact position is determined based on the user-defined convergence factor: $messBW_{i+1}=messBW_{i}+convFactor\times(cpuBW_{i}-messBW_{i})$. The approach is based on the proportional–integral controller mechanism from the control theory [77, 78]. Finally, the Mess uses $messBW_{i+1}$ to read the $Latency_{i+1}$ from the corresponding bandwidth–latency memory curve. The next simulation window starts with the Mess simulator providing memory access $Latency_{i+1}$ to the CPU simulator.
@@
Figure 7: Mess simulator uses a feedback control loop to adjust the simulated memory access latency.
@@
III-B Evaluation
We evaluate the Mess simulator integrated with ZSim and gem5 against the actual hardware. We compare the simulated and actual bandwidth–latency curves as well as the performance of memory-bound benchmarks: STREAM [5], LMbench [8], and Google multichase [9].
III-B1 ZSim
Figure 9 shows the DDR4, DDR5 and HBM2 bandwidth–latency curves measured with the ZSim connected to the Mess simulator. [^5] Mess simulator also supports the Intel Optane technology. Optane’s bandwidth–latency curves are measured on a 16-core Cascade Lake server with 6$\times$×DDR4-2666 16 GB and 2$\times$×Intel Optane 128 GB memory in App Direct mode. Since Intel Optane technology is discontinued since 2023, we do not analyze its performance characteristics and simulation. The configurations of the simulators match the actual Intel Skylake with 24-core and six DDR4-2666 memory channels. The simulated Mess curves, depicted in Figure 9(a), closely resemble the actual memory systems behavior (Figure 3(a)). The simulation error of the unloaded memory latency is below 1%, and it is around 3% for the maximum latencies. The difference between the simulated and the actual saturated bandwidth range is only 2%. Figures 9(b) and 9(c) show the ZSim+Mess simulation results for the high-end DDR5 and HBM memories. To saturate the 8-channel DDR5-4800 and 32-channel HBM2, we increase the number of simulated cores to 58 and 192, respectively. Again, the simulated bandwidth–latency curves closely resemble the performance of the corresponding actual memory systems (Figures 3(g) and 3(e)).
@@ III-B1 ZSim
III-B2 gem5
Figure 11 shows the DDR5 and HBM2 bandwidth–latency curves simulated with the gem5 connected to the Mess memory simulator. In all experiments, the gem5 is configured to model Graviton 3 cores [59] described in Section II-E1. To reduce simulation time, we simulate 16 CPU cores connected to a single memory channel. [^6] Simulation of a whole server comprising 64 Graviton 3 cores and 8$\times$×DDR5-4800 requires more than five hours to obtain a single bandwidth–latency datapoint. The full simulation of the bandwidth–latency curves would require more than a year. The simulated bandwidth–latency curves, when scaled to eight DDR5 channels or 32 HBM2 channels, closely resemble actual system behavior (Figures 3(e) and 3(g)).
We also evaluate the Mess memory simulation against the gem5’s built-in simple memory model and internal DDR5 model when running STREAM, LMbench, and Google Multichase benchmarks. In these experiments we simulate a whole server comprising 64 Graviton 3 cores and 8$\times$×DDR5-4800, and compare the results against the benchmark executions on the actual server. The results are presented in Figure 12. The Mess simulator decreases the average error from 30% (gem5 simple memory model) and 15% (internal DDR5 model) to only 3%. Such a low error is unprecedented in any prior validation attempts [79, 80, 81, 82]. Moreover, the gem5 with Mess simulator reduces the simulation time by 1.6% w.r.t. to much simpler built-in memory models.
@@
III-C Simulation of novel memory systems: CXL memory expanders
The Mess simulator provides a fundamental solution for this gap because it can simulate emerging memory systems as soon as their bandwidth–latency curves are available. For memory technologies available on the market the curves can be measured on a real platform. For emerging memory devices that are not yet available in off-the-shelf servers, the bandwidth–latency curve can be measured on a developer board with a prototype of the new device, or alternatively it can be provided by the manufacturers, e.g. based on their detailed proprietary RTL models.
The Mess simulation is performed with the CXL memory expanders bandwidth–latency curves provided by Micron Technology based on their detailed hardware model. In particular, we model a CXL memory expander connected to the host via the CXL 2.0 PCIe 5.0 interface with 1$\times$8 Lanes. The device comprises one memory controller connected to a DDR5-5600 DIMM with two ranks. All the CXL modules, Front end, Central controller and Memory controller, are implemented in SystemC. The modules communicate by using the manufacturer’s proprietary Transaction Level Modeling (TLM) framework, which is based on SystemC TLM [83].
The obtained bandwidth–latency curves are shown in Figure 13(a). The figure plots the round-trip latency from the CXL host input pins. To consider a full load-to-use latency, a user should add the round-trip time between the CPU core and the CXL host. We measure this latency component with the Intel MLC [11]. The CXL memory expanders show a similar performance trends as the DDRx or HBM memory systems: latency that increases with the system load, significant non-linear increase after a saturation point, and the impact of the traffic read/write ratio. One major difference is that the CXL interface provides the best performance for a balanced reads and writes traffic, while its performance drops significantly for the 100%-read or 100%-write traffic. This is because, unlike the DDRx interfaces, CXL is a full-duplex interconnect with independent read and write links. Therefore, the CXL can transmit simultaneously in both directions, but in the case of the unbalanced traffic one CXL transmission direction could be saturated while other direction is negligibly used.
We use the obtained CXL memory bandwidth–latency curves in the Mess simulator integrated with ZSim, gem5 and OpenPiton Metro-MPI simulators (see Figure 13). To reduce long OpenPiton Metro-MPI simulation time, we model only 25 curves with a small number of experimental points in each curve. For this reason some segments of the curves are discrete. Nevertheless, the OpenPiton Metro-MPI simulations match the general trend and the saturated bandwidth range of the manufacturer’s curves. The maximum latency range is below the manufacturers CXL curves because the simulated small in-order Ariane cores with only 2-entry MSHRs cannot saturate the target memory system. This behavior is already detected and discussed in Section II-E3. ZSim and gem5 results practically match the manufacturer’s CXL curves. In Appendix -B, we compare our CXL simulation platform against prior approach to emulate memory-over-CXL.
@@ III-C Simulation of novel memory systems: CXL memory expanders
IV Mess application profiling
@@
IV-A Background: Extrae and Paraver
@@
Extrae automatically collects entry and exit call points to the programming model runtime, source code references, hardware counters metrics, dynamic memory allocation, I/O system calls, and user functions. It is is compatible with programs written in C, Fortran, Java, Python, and combinations of different languages, and it supports a wide range of parallel programming models. Extrae is available for most UNIX-based operating systems and it is deployed in all relevant HPC architectures, including CPU-based systems and accelerators.
IV-B Use cases
@@
IV-B1 Bandwidth–latency curves
Figure 14 depicts the Mess memory-related profile of the HPCG benchmark. Most of the HPCG execution is located in the saturated bandwidth area, above 75 GB/s. Sporadically, the benchmark even reaches the maximum sustained bandwidth with peak memory latencies in the range of 260–290 ns. Also, each HPCG point on the curves is associated with a memory stress score. The score value ranges from 0, for the unloaded memory system, to 1, corresponding to the right-most area of the bandwidth–latency curves. Memory stress score in a given point is calculated as a weighted sum of two parameters, the memory latency and the curve inclination. The latency itself is a good proxy of the system stress, while the inclination shows the memory system sensitivity to a bandwidth change. Gentle inclination indicates that a memory bandwidth change would have a minor impact on the memory access latency and the overall performance. In the steep curve segments, e.g. 95–100 GB/s area in Figure 14, small bandwidth changes can rapidly saturate the memory system leading to a major latency increase. The Mess extension of Paraver already includes the stress score visualization with a green–yellow–red gradient that can be easily interpreted by application developers.
IV-B2 Timeline analysis
Once the memory stress score is incorporated into the application’s Paraver trace, it can be combined with other aspects of the application analysis, as illustrated in Figure 15. The figure analyzes around two seconds of the HPCG runtime, from 241,748,818 $\mu$$s$ to 243,728,242 $\mu$$s$ ($x$-axis). Guided by the sequence of MPI calls illustrated in Figure 15 (top), we identify the application’s main iterative loop and, using MPI_Allreduce (pink) as delimiter, we select two iterations to focus the analysis on the compute phases (Figure 15, middle). The color gradient corresponds to the length of the compute phase: green to blue gradient for short to long phases. Figure 15 (bottom) shows the memory stress score for this region. The longest compute phases (blue) exhibits two distinct memory pressure behaviors: at the start of the phase, the memory stress score rises to 0.71, and then halfway through the phase it decreases to 0.64. The fine-grain application profiling can detect different memory stress score values even within a single compute phase.
@@ IV-B2 Timeline analysis
IV-B3 Links to the source code
@@ V Related work
V-A Memory system benchmarks
@@
V-B Memory system simulation
@@
The Mess framework tightly integrates the memory performance characterization into the memory simulation. We compare the Mess simulator with the state-of-the-art memory models included in the CPU simulators [14, 13, 18, 75, 74, 73, 80, 81, 46] as well as the external cycle-accurate memory simulators [22, 24]. The Mess memory simulator is fast, accurate and easily-integrated with the CPU simulators. It can support novel memory systems as soon as their detailed bandwidth–latency curves are measured on actual production systems or prototypes, or provided by the manufacturers. Therefore, Mess can simulate high-end and future memory systems much sooner than the standard memory simulators which consider detailed memory device sequences and timings [21, 24, 22, 23, 25, 20]. Mess is the first memory simulator to support CXL memory expanders.
@@
Apart from the hardware simulation, system performance can be analyzed with analytical models. The PROFET model [53] and Interval model [93] predict memory system impact on the overall application performance. The predictions are based on the application runtime profiling combined with an analytical performance model. The main objective is to provide an alternative to complex and slow CPU simulations. This is orthogonal and complementary to the Mess simulator that targets the main memory simulation.
V-C Application profiling

### v4

@@ V Mess simulator
In this section we will present the Mess analytical memory system simulator and show how it significantly improves the memory simulation accuracy and enables a quick adoption of new memory technologies in hardware simulators.
V-A Design
@@
The CPU and memory simulators are typically connected in the following way: the CPU simulator issues the memory operations and the memory simulator determines their latencies. The Mess simulator does this analytically based on the application’s position in the memory bandwidth–latency curves. This process is complex due to the inherent dependency between the memory system latency, timings of the memory operations and all dependent instructions, and the generated memory bandwidth. We simplify the problem by designing the Mess simulator not to compute the exact memory latency for a given memory traffic, but to detect and correct discrepancies between the memory access latency and the simulated bandwidth. This approach, together with the fundamental principle of application’s position in the memory bandwidth–latency curves, enables the Mess simulator to surpass the accuracy of all other memory simulators, while remaining simple and fast. The only memory system parameter required by the Mess model is the family of the bandwidth–latency curves. The curves can be measured on the actual hardware (Section III) or can be provided by the manufacturers, e.g. based on their detailed hardware model, as we will discuss in Section V-C.
@@
The Mess simulator acts as a feedback controller [83] from the classical control theory [84], illustrated in Figure 8. The simulation can start from any memory access latency, e.g. the unloaded one. This latency is used by the CPU simulator which generates memory memory reads and writes. The Mess simulator observers this simulated memory bandwidth, positions it at the corresponding memory bandwidth–latency curve, and controls whether it coincides with the memory latency used in the CPU simulation. If this is not the case, the memory latency is being adjusted with an iterative process we describe later.
The control process is performed at the end of each simulation window, which, in our experiments, comprises 1000 memory operations. This is much smaller than the length of the application phases [66], so the transition error between different application phases has a negligible impact on the simulator’s accuracy.
@@
Figure 8: Mess feedback control loop adjusts the simulated memory access latency.
@@
Figure 9 describes one iteration of the Mess simulator control loop. For simplicity, the figure shows a single bandwidth–latency curve. We start with the Mess estimate of the application’s bandwidth–latency position in the $i^{th}$ simulation window, $(messBW_{i},~{}Latency_{i})$1. The curve is selected based on the read/write ratio of the simulated memory traffic, and the $x$-axis position is determined by $messBW_{i}$. From that point on, all the issued memory requests are simulated with $Latency_{i}$. At the end of the simulation window, the Mess simulator monitors the simulated memory bandwidth, $cpuBW_{i}$2, and compares it with the $messBW_{i}$ estimated at the beginning of the window  3.
If the simulation is in a steady-state and the application did not change its behavior, there will be no major difference between $cpuBW_{i}$ and $messBW_{i}$. This confirms the consistency in the simulated memory access latency, the CPU timings and the achieved memory bandwidth. Therefore, the CPU simulation in the next window will continue with the same memory latency.
Otherwise, a difference between $cpuBW_{i}$ and $messBW_{i}$ suggests inconsistent simulated memory latency and bandwidth. This can happen, for example, if the application changes its behavior. Figure 9 illustrates the case in which the application increases the frequency of memory request leading to the higher bandwidth: $cpuBW_{i}>messBW_{i}$3. In this case, the simulated memory bandwidth $cpuBW_{i}$ does not correspond to the memory $Latency_{i}$ used in the CPU simulation. To address this inconsistency, the Mess simulator adjusts the predicted application position in the bandwidth–latency curves. The objective of this adjustment is not to reach the correct $(BW,Latency)$ position in a single iteration. The next Mess estimate, $(messBW_{i+1},~{}Latency_{i+1})$, will be positioned in-between $messBW_{i}$ and $cpuBW_{i}$4. The exact position is determined based on the user-defined convergence factor: $messBW_{i+1}=messBW_{i}+convFactor\times(cpuBW_{i}-messBW_{i})$. The approach is based on the proportional–integral controller mechanism from the control theory [83, 84]. Finally, the Mess uses $messBW_{i+1}$ to read the $Latency_{i+1}$ from the corresponding bandwidth–latency curve. The $Latency_{i+1}$ is a load-to-use latency. This includes the time spent in the CPU cores, cache hierarchy and network on chip, already considered in the CPU simulation. In the final step, the $Latency_{i+1}$ is adjusted by these CPU timings $Latency_{i+1}^{Memory}=Latency_{i+1}-Latency_{i+1}^{CPU}$. The next simulation window starts with the Mess providing the updated $Latency_{i+1}^{Memory}$ to the CPU simulator.
V-B Evaluation
The Mess simulator is integrated with ZSim and gem5, and evaluated against the actual hardware. We compare the simulated and actual bandwidth–latency curves as well as the performance of memory-bound benchmarks: STREAM [5], LMbench [8], and Google multichase [9].
V-B1 ZSim
Figure 10 shows the DDR4, DDR5 and HBM2 bandwidth–latency curves measured with the ZSim connected to the Mess simulator. [^3] The Mess simulator also supports the Intel Optane technology. Optane’s bandwidth–latency curves are measured on a 16-core Cascade Lake server with 6$\times$×DDR4-2666 16 GB and 2$\times$×Intel Optane 128 GB memory in App Direct mode. Intel Optane technology is discontinued since 2023, so we do not analyze its performance characteristics and simulation. The configurations of the simulators match the actual Intel Skylake with 24-core and six DDR4-2666 memory channels. The simulated Mess curves, depicted in Figure 10(a), closely resemble the actual memory systems performance (Figure 3(a)). The simulation error of the unloaded memory latency is below 1%, and it is around 3% for the maximum latencies. The difference between the simulated and the actual saturated bandwidth range is only 2%. Figures 10(b) and 10(c) show the ZSim+Mess simulation results for the high-end DDR5 and HBM memories. To saturate the 8-channel DDR5-4800 and 32-channel HBM2, we increase the number of simulated cores to 58 and 192, respectively. Again, the simulated bandwidth–latency curves closely resemble the performance of the corresponding actual memory systems (Figures 3(g) and 3(e)).
@@ V-B1 ZSim
V-B2 gem5
Figure 12 shows the DDR5 and HBM2 bandwidth–latency curves simulated with the gem5 connected to the Mess memory simulator. In all experiments, the gem5 is configured to model Graviton 3 cores [61] described in Section IV-A. To reduce simulation time, we simulate 16 CPU cores connected to a single memory channel. [^4] Simulation of a whole server comprising 64 Graviton 3 cores and 8$\times$×DDR5-4800 requires more than five hours to obtain a single bandwidth–latency datapoint. The full simulation of all the curves would require more than a year. The simulated bandwidth–latency curves, when scaled to eight DDR5 channels or 32 HBM2 channels, closely resemble actual system behavior (Figures 3(e) and 3(g)).
We also evaluate the Mess memory simulation against the gem5’s built-in simple memory model and internal DDR5 model as well as cycle-accurate Ramulator2 when running STREAM, LMbench, and Google Multichase benchmarks. In these experiments we simulate a whole server comprising 64 Graviton 3 cores and 8$\times$×DDR5-4800, and compare the results against the benchmark executions on the actual server. The gem5+Mess simulation time is practically the same as the gem5 internal DDR model, while it provides much better accuracy. The Mess memory simulator decreases the average error from 30% (gem5 simple memory model), 15% (internal DDR5 model), and 52% (Ramulator 2) to only 3%. Such a low error is unprecedented in any prior validation attempts [85, 86, 87, 88].
V-C Simulation of novel memory systems: CXL memory expanders
The Mess simulator provides a fundamental solution for this gap because it can simulate emerging memory systems as soon as their bandwidth–latency curves are available. For memory technologies available on the market, the curves can be measured on a real platform. For emerging memory devices that are not yet available in off-the-shelf servers, the bandwidth–latency curve can be measured on a developer board with a prototype of the new device, or alternatively it can be provided by the manufacturers, e.g. based on their detailed proprietary RTL models.
The Mess simulation is performed with the CXL memory expanders bandwidth–latency curves provided by the memory manufacturer based on their detailed hardware model. In particular, we model a CXL memory expander connected to the host via the CXL 2.0 PCIe 5.0 interface with 1$\times$8 Lanes. The device comprises one memory controller connected to a DDR5-5600 DIMM with two ranks. All the CXL modules, Front end, Central controller and Memory controller, are implemented in SystemC. The modules communicate by using the manufacturer’s proprietary SystemC Transaction Level Modeling (TLM [89]) framework,
@@ V-C Simulation of novel memory systems: CXL memory expanders
The obtained bandwidth–latency curves are shown in Figure 14(a). [^5] The maximum theoretical bandwidth of the CXL.mem protocol is influenced by the read/write ratio of the workload [90]. In this figure, we present the highest value among all possible scenarios. The figure plots the round-trip latency from the CXL host input pins. To consider a full load-to-use latency, a user should add the round-trip time between the CPU core and the CXL host. We measure this latency component with the Intel MLC [11]. The CXL memory expanders show a similar performance trends as the DDRx or HBM memory systems: latency that increases with the system load, significant non-linear increase after a saturation point, and the impact of the traffic read/write ratio. One major difference is that the CXL interface provides the best performance for a balanced reads and writes traffic, while its performance drops significantly for the 100%-read or 100%-write traffic. This is because, unlike the DDRx interfaces, CXL is a full-duplex interconnect with independent read and write links. Therefore, the CXL can transmit simultaneously in both directions, but in the case of the unbalanced traffic one CXL transmission direction could be saturated while other direction is negligibly used.
We use the obtained CXL memory bandwidth–latency curves in the Mess simulator integrated with ZSim, gem5 and OpenPiton Metro-MPI (see Figure 14). In all configurations, the Mess simulator closely follows the Manufacturer’s SystemC model. To reduce long OpenPiton Metro-MPI simulation time, we model only 25 curves with a small number of experimental points in each curve. For this reason some segments of the curves are discrete. Nevertheless, the OpenPiton Metro-MPI simulations match the general trend and the saturated bandwidth range of the manufacturer’s curves. The maximum latency range is below the manufacturers CXL curves because the simulated small in-order Ariane cores with only 2-entry MSHRs cannot saturate the target memory system. This behavior is already detected and discussed in Section IV-C. ZSim and gem5 results practically match the manufacturer’s CXL curves. In Appendix -B, we compare our CXL simulation platform against prior approach to emulate memory-over-CXL.
VI Mess application profiling
@@
VI-A Background: Extrae and Paraver
@@
Extrae automatically collects entry and exit call points to the programming model runtime, source code references, hardware counters metrics, dynamic memory allocation, I/O system calls, and user functions. It is is compatible with programs written in C, Fortran, Java, Python, and combinations of different languages. It supports a wide range of parallel programming models. Extrae is available for most UNIX-based operating systems and it is deployed in all relevant HPC architectures, including CPU-based systems and accelerators.
VI-B Use cases
@@
VI-B1 Bandwidth–latency curves
Figure 15 depicts the Mess profile of the HPCG benchmark. Most of the HPCG execution is located in the saturated bandwidth area, above 75 GB/s. Sporadically, the benchmark even reaches the maximum sustained bandwidth with peak memory latencies in the range of 260–290 ns. Also, each HPCG point on the curves is associated with a memory stress score. The score value ranges from 0, for the unloaded memory system, to 1, corresponding to the right-most area of the bandwidth–latency curves. Memory stress score in a given point is calculated as a weighted sum of the memory latency and the curve inclination. The latency itself is a good proxy of the system stress, while the inclination shows the memory system sensitivity to a bandwidth change. Gentle inclination indicates that a memory bandwidth change would have a minor impact on the memory access latency and the overall performance. In the steep curve segments, e.g. 95–100 GB/s area in Figure 15, small bandwidth changes can rapidly saturate the memory system leading to a major latency increase. The Mess extension of Paraver already includes the stress score visualization with a green–yellow–red gradient that can be easily interpreted by application developers.
VI-B2 Timeline analysis
Once the memory stress score is incorporated into the application’s Paraver trace, it can be combined with other aspects of the application analysis, as illustrated in Figure 16. The figure analyzes around two seconds of the HPCG runtime, from 241,748,818 $\mu$$s$ to 243,728,242 $\mu$$s$ ($x$-axis). Guided by the sequence of MPI calls illustrated in Figure 16 (top), we identify the application’s main iterative loop and, using MPI_Allreduce (pink) as delimiter, we select two iterations for our analysis. The middle Figure 16 analyzes compute applications phases. The color gradient corresponds to the compute phase length: green to blue gradient for short to long phases. Figure 16 (bottom) shows the memory stress score for this region. The longest compute phases (blue) exhibits two distinct memory behaviors: at the start of the phase, the memory stress score rises to 0.71, and then halfway through the phase it decreases to 0.64. The fine-grain application profiling can detect different memory stress score values even within a single compute phase.
VI-B3 Links to the source code
@@ VII Related work
VII-A Memory system benchmarks
@@
VII-B Memory system simulation
@@
The Mess framework tightly integrates the memory performance characterization into the memory simulation. We compare the Mess analytical memory system simulator with the state-of-the-art memory models included in the CPU simulators [14, 13, 18, 78, 77, 76, 86, 87, 49] as well as the external cycle-accurate memory simulators [22, 23, 24]. The Mess memory simulator is fast, accurate and easily-integrated with the CPU simulators. It can support novel memory systems as soon as their detailed bandwidth–latency curves are measured on actual production systems or prototypes, or provided by the manufacturers. Therefore, Mess can simulate high-end and future memory systems much sooner than the standard memory simulators which consider detailed memory device sequences and timings [21, 23, 22, 24, 25, 20]. Mess is the first memory simulator to support CXL memory expanders.
@@
Apart from the hardware simulation, system performance can be analyzed with analytical models. The PROFET model [55] predicts how an application’s performance and energy consumption change when it is executed on different (future) memory systems. The model is based on the instrumentation of an application execution on actual hardware, so it already takes into account CPU microarchitectural details. PROFET is evaluated on Intel and Huawei servers with DDR3, DDR4, HBM and Optane memory [101]. The main objective is to provide an alternative to complex and slow CPU simulations. This is orthogonal and complementary to the Mess simulator that targets the main memory simulation.
VII-C Application profiling